Integrate the following with respect to x :

(a) (5x-3)^4
1
Expert's answer
2020-06-21T16:45:05-0400

We can integrate the function "(5x-3)^4" by substitution. Set "u=5x-3", then "du=5dx," and "dx=\frac{du}{5}."

Hence

"\int (5x-3)^4dx=\int u^4\frac{du}{5}=\frac{1}{5}\int u^4du=\frac{1}{5}\cdot \frac{u^5}{5}+C=\frac{u^5}{25}+C="

"=\frac{(5x-3)^5}{25}+C."